# Content-Oblivious Leader Election in 2-Edge-Connected Networks

### v2

@@ 2.2 Leader Election
More precisely, each node maintains a binary internal variable $\mathsf{isLeader}$ to represent its current output, taking the value $\mathsf{Leader}$ or $\mathsf{Non\text{-}Leader}$. The value of $\mathsf{isLeader}$ may change arbitrarily many times during the execution of the algorithm, but it must eventually stabilize to a final value that remains unchanged thereafter. The final outputs of all nodes must collectively satisfy the conditions of Definition 2.
@@
Ideally, we aim to design leader election algorithms that are quiescently terminating, as this property allows a second algorithm to be executed safely after leader election. In contrast, if the algorithm guarantees only stabilization or termination, nodes may be unable to distinguish whether an incoming pulse belongs to the leader election process or the subsequent algorithm. This ambiguity can affect the correctness of both algorithms. For this reason, prior work [content-oblivious-leader-election-24] establishes a general algorithm simulation result only for oriented rings, but not for unoriented rings.
@@ 2.3 DFS and Strongly Connected Orientation
We describe a DFS tree $T_{G}$ of the graph $G=(V,E)$ that is uniquely determined by the port numbering and node identifiers. Based on this DFS tree, we describe an edge orientation of $G$ that transforms it into a strongly connected directed graph $\vec{G}$. Both the DFS tree and the orientation play a critical role in the analysis of our algorithm.
@@
$\vec{E}_{T}$ =$=\{(u,v)\,|\,\{u,v\}\in E_{T}\text{ and }u\text{ is the parent of}v\text{ in }T_{G}\},$
$\vec{E}_{B}$ =$=\{(u,v)\,|\,\{u,v\}\in E_{B}\text{ and }v\text{ is an ancestor of }u\text{ in }T_{G}\}.$
A directed graph is strongly connected if for any two vertices $u$ and $v$ in the graph, there is a directed path from $u$ to $v$. We have the following observation.
Observation 8.
@@
To prove the observation, it suffices to show that, for any tree edge $e=\{u,v\}$, both $u$ and $v$ are in the same strongly connected component of $\vec{G}$, as this implies that all nodes in $V$ are in the same strongly connected component. Let $u$ be the parent of $v$. By Definition 7, we already know that $(u,v)$ is a directed edge of $\vec{G}$, so it remains to show that $u$ is reachable from $v$ in $\vec{G}$.
@@
Throughout the paper, for any two nodes $u\in V$ and $v\in V$, we write $P_{u,v}$ to denote any shortest path from $u$ to $v$ in $\vec{G}$, whose existence is guaranteed by 8. We allow $u=v$, in which case $P_{u,v}$ is path of zero length consisting of a single node $u=v$.
@@ 3 Technical Overview
[…]
3.1 Synchronized Counting
@@
[…]
@@
After electing itself as the leader, node $v$ freezes its counter at $\mathsf{Count}(v)=\mathsf{ID}(v)$ and does not increment it further. This action caps the counter value of any neighbor of $v$ at most $\mathsf{ID}(v)+1$. More generally, in a graph of $n\leq N$ nodes, the counter value of any node is upper bounded by $\mathsf{ID}(v)+N-1$, since any two nodes are connected by a path of length at most $N-1$.
@@ 3.2 DFS Notification
Intuitively, the leader $r$ is able to notify a neighbor $u$ because the counter value of node $u$ is constrained by a chain $P$, whose other end is anchored at the leader $r$, whose counter has already been frozen. To extend this notification process to the entire network, we perform a DFS, using the same notification procedure to explore each new edge. When a node finishes exploring all its remaining incident edges, it can notify its parent by sending pulses until a total of $\mathsf{ID}(r)+N+2$ pulses have been sent through this port since the start of the algorithm. When a node $u$ explores a back edge $\{u,v\}$, the node $v$ replies using the same method. When a node decides which new incident edge to explore, it prioritizes the edge with the smallest port number, so the resulting DFS tree is $T_{G}$, as defined in Section 2.3.
@@
Recall from Section 2.3 that $P_{a,b}$ is a shortest path from $a$ to $b$ in the directed version $\vec{G}$ of $G$ defined in Definition 7.
@@
Both our leader election algorithm and the simulation algorithm of Censor-Hillel, Cohen, Gelles, and Sela [fully-defective-22] utilize DFS, but for different purposes. They perform a DFS from a pre-selected leader to find a cycle, while we use DFS to disseminate the identity of the elected leader to all nodes and to achieve quiescent termination. Coincidentally, both approaches leverage the strongly connected orientation of 2-edge-connected graphs. In our case, the orientation $\vec{G}$ is used solely in the analysis to facilitate the chain-and-anchor argument. In their case, it guarantees the existence of a Robbins cycle, which plays a critical role in their algorithm.
4 Quiescently Terminating Leader Election
@@
The algorithm for Theorem 1 is presented in Algorithm 1, with the DFS notification subroutine detailed in Algorithm 2. There are five main variables used in our algorithm:
@@
For ease of analysis, several positions in Algorithm 1 and Algorithm 2 are annotated with labels: $\mathsf{StartDFS}(v)$ (node $v$ initiates the DFS), $\mathsf{SendExplore}_{i}(v)$ (node $v$ begins sending an explore-notification on port $i$), $\mathsf{ReceiveExplore}_{i}(v)$ (node $v$ confirms receipt of an explore-notification from port $i$), $\mathsf{SendDone}_{i}(v)$ (node $v$ begins sending a done-notification on port $i$), and $\mathsf{ReceiveDone}_{i}(v)$ (node $v$ confirms receipt of a done-notification from port $i$). These labels are treated as events in the analysis.
@@
Any pulse that is sent during Algorithm 2 is called a DFS notification pulse. Any remaining pulse is called a synchronized counting pulse. This distinction is used only in the analysis. The nodes cannot see these labels in the algorithm.
@@
  // Forcing all $\mathsf{ID}$s to be integer multiples of $N$
@@ 19                   $\mathsf{Leader}\mathsf{ID}\leftarrow\lfloor\mathsf{Count}/N\rfloor\cdot N$
21                   Event: $\mathsf{ReceiveExplore}_{j}(v)$
@@ 13                         Event: $\mathsf{ReceiveDone}_{j}(v)$
15                   else  // Incoming DFS from port $h$ via an back edge
@@ 4.1 Synchronized Counting
We begin with analyzing the synchronized counting phase of our algorithm, which corresponds to the main loop of Algorithm 1. We aim to show that the node $r$ with the smallest identifier eventually elects itself as the leader ($\mathsf{State}=\mathsf{Leader}$) while all other nodes are still in the synchronized counting phase ($\mathsf{State}=\bot$).
Algorithm 1 starts with $\mathsf{Count}=0$. The action $\mathsf{SendAll}(1)$ is performed upon initialization and immediately after each counter increment, so $\sigma_{i}(v)=\mathsf{Count}(v)+1$ for all $i\in[\deg(v)]$. The counter is incremented by one whenever an incoming pulse causes $\min_{j\in[\deg(v)]}\rho_{j}$ to increase, so we always have $\mathsf{Count}=\min_{j\in[\deg(v)]}\rho_{j}$. There are two ways to exit the main loop of Algorithm 1. The first is $\mathsf{Count}=\mathsf{ID}(v)$, which causes $v$ to become a leader. The second is $\rho_{i}-\sigma_{i}>1$, which causes $v$ to become a non-leader.
Within Section 4.1, we focus on analyzing the behavior of the nodes in the synchronized counting phase. The only fact that we need to know about the DFS notification phase is that upon exiting the main loop of Algorithm 1, the value of $\mathsf{Count}$ is permanently fixed. The analysis of the DFS notification phase is done in Section 4.2.
Observation 9 (Outgoing synchronized counting pulses).
@@
During the synchronized counting phase, the action $\mathsf{SendAll}(1)$ is executed during initialization and immediately after each counter increment. There is no other mechanism by which node $v$ sends synchronized counting pulses in the algorithm. Since $\mathsf{Count}(v)$ starts at zero and becomes fixed when $v$ exits the main loop of Algorithm 1, it follows that for any node $v\in V$, at any time during the algorithm, the number of synchronized counting pulses sent on each of its ports equals $\mathsf{Count}(v)+1$. ∎
@@
Observation 10 (Counter value).
@@
Since $\mathsf{Count}(v)$ starts at zero, it suffices to show that, while $v$ is in the synchronized counting phase, the counter is incremented by one whenever an incoming pulse causes $\min_{j\in[\deg(v)]}\rho_{j}$ to increase, and that there is no other mechanism for changing the counter value. This follows from the observation that a pulse arriving on port $i$ increases $\min_{j\in[\deg(v)]}\rho_{j}$ if and only if $\rho_{i}=\min_{j\in[\deg(v)]}\rho_{j}$ immediately after the pulse arrives. This is precisely the condition for incrementing the counter in Algorithm 1. ∎
Since the counter is frozen once a node exits the synchronized counting phase, there is a simple generalization of 10 to any node $v$ in any phase: $\mathsf{Count}(v)\leq\min_{j\in[\deg(v)]}\rho_{j}(v)$.
@@
Suppose the port number of $e$ is $j$ at node $u$ and $i$ at node $v$. Since the counter is frozen once a node exits the synchronized counting phase, 10 implies that $\mathsf{Count}(u)\leq\rho_{j}(u)$. In other words, the number of pulses from $v$ that have been delivered to $u$ is at least $\mathsf{Count}(u)$. 9, together with the assumption that no DFS notification pulse from $v$ has been delivered to $u$, implies that $u$ receives at most $\mathsf{Count}(v)+1$ pulses from $v$, so $\rho_{j}(u)\leq\mathsf{Count}(v)+1$. Combining the two inequalities, we obtain $\mathsf{Count}(u)\leq\mathsf{Count}(v)+1$. Similarly, we also have $\mathsf{Count}(v)\leq\mathsf{Count}(u)+1$, so $|\mathsf{Count}(u)-\mathsf{Count}(v)|\leq 1$. ∎
@@
Let $u$ be the neighbor of $v$ such that the port number of $e=\{u,v\}$ at node $v$ is $i$. While $v$ is in the synchronized counting phase, by 9, $\sigma_{i}(v)=\mathsf{Count}(v)+1$. If $v$ has not received any DFS notification pulse from $u$, then 9 and Lemma 11 imply that the number of pulses $v$ received from $u$ is $\rho_{i}(v)\leq\mathsf{Count}(u)+1\leq\mathsf{Count}(v)+2$, so
@@
For the rest of the discussion, we prove the claim. For each neighbor $u$ of $v$ in $S$, we have $\mathsf{Count}(u)\geq c$. By 9, $u$ has already sent $\mathsf{Count}(u)+1\geq c+1$ pulses to $v$. If (C1) has not happened yet, for each neighbor $w$ of $v$ in $V\setminus S$, we also have $\mathsf{Count}(w)\geq c$. By 9, $w$ has already sent $\mathsf{Count}(w)+1\geq c+1$ pulses to $v$.
@@
If, by this point, the $\mathsf{Non\text{-}Leader}$ exit condition $\rho_{i}(v)-\sigma_{i}(v)>1$ for $v$ has not been triggered yet, by 10, $\mathsf{Count}(v)$ will be incremented to $c+1$, as required.
Now suppose the $\mathsf{Non\text{-}Leader}$ exit condition $\rho_{i}(v)-\sigma_{i}(v)>1$ is triggered due to some incoming pulse at some point. By Lemma 12, the pulse must be a DFS notification pulse. If this pulse comes from a node in $V\setminus S$, then (C3) has happened. Otherwise, the pulse comes from a node in $S$, meaning that this node has already exited the synchronized counting phase, so (C2) has happened. ∎
@@
By Lemma 12, the first node to exit the synchronized counting phase cannot do so via the $\mathsf{Non\text{-}Leader}$ exit condition. Hence it suffices to show that $r$ is the first node that exits the synchronized counting phase.
Apply Lemma 13 with $S=V$. Since $V\setminus S=\emptyset$, (C3) is vacuously false. Moreover, since $\min_{u\in V\setminus S}\mathsf{Count}(u)=\infty$, (C1) reduces to $\min_{v\in S}\mathsf{Count}(v)=\infty$, which cannot hold: By the time the first node $v$ reaches $\mathsf{Count}(v)=\mathsf{ID}(v)$, it must have already exited the synchronized counting phase and frozen its counter. Therefore, we conclude that (C2) must eventually occur, and hence some node $v$ eventually exits the synchronized counting phase.
@@
Suppose $v\neq r$. Consider the first moment when $\mathsf{Count}(v)=\mathsf{ID}(v)$. Consider any path $P$ connecting $v$ and $r$. Since no DFS notification pulse has been delivered so far, by Lemma 11, the counter values of any two adjacent nodes in $P$ differ by at most one, so
@@
If we simply let the elected leader $r$ halt after the leader election phase and do not execute Algorithm 2, we obtain a simple stabilizing leader election algorithm. By Lemma 11, for any node $v\in V\setminus\{r\}$, we have $\mathsf{Count}(v)\leq\mathsf{Count}(r)+\mathsf{length}(P)<\mathsf{ID}(r)+(N-1)<\mathsf{ID}(v)$, so the $\mathsf{Leader}$ exit condition cannot be triggered.

### v3

@@ 2.2 Leader Election
More precisely, each node maintains a binary internal variable $\mathsf{isLeader}$ to represent its current output, taking the value $\mathsf{Leader}$ or $\mathsf{Non\text{-}Leader}$. The value of $\mathsf{isLeader}$ may change arbitrarily many times during the execution of the algorithm, but it must eventually stabilize to a final value that remains unchanged thereafter. The final outputs of all nodes must collectively satisfy the conditions of Definition˜3.
@@
Ideally, we aim to design leader election algorithms that are quiescently terminating, as this property allows a second algorithm to be executed safely after leader election. In contrast, if the algorithm guarantees only stabilization or termination, nodes may be unable to distinguish whether an incoming pulse belongs to the leader election process or the subsequent algorithm. This ambiguity can affect the correctness of both algorithms. For this reason, prior work [FGG+24] establishes a general algorithm simulation result only for oriented rings, but not for unoriented rings.
@@ 2.3 DFS and Strongly Connected Orientation
We describe a DFS tree $T_{G}$ of the graph $G=(V,E)$ that is uniquely determined by the port numbering and node identifiers. Based on this DFS tree, we describe an edge orientation of $G$ that transforms it into a strongly connected directed graph $\vec{G}$. Both the DFS tree and the orientation play a critical role in the analysis of our algorithm for Theorem˜1.
@@
[…]
A directed graph is strongly connected if for any two nodes $u$ and $v$ in the graph, there is a directed path from $u$ to $v$. We have the following observation.
Observation 9.
@@
To prove the observation, it suffices to show that, for any tree edge $e=\{u,v\}$, both $u$ and $v$ are in the same strongly connected component of $\vec{G}$, as this implies that all nodes in $V$ are in the same strongly connected component. Let $u$ be the parent of $v$. By Definition˜8, we already know that $(u,v)$ is a directed edge of $\vec{G}$, so it remains to show that $u$ is reachable from $v$ in $\vec{G}$.
@@
Throughout the paper, for any two nodes $u\in V$ and $v\in V$, we write $P_{u,v}$ to denote any shortest path from $u$ to $v$ in $\vec{G}$, whose existence is guaranteed by ˜9. We allow $u=v$, in which case $P_{u,v}$ is a path of zero length consisting of a single node $u=v$.
@@ 3 Technical Overview
We now give a high-level overview of the proofs for our Theorems˜1 and 2. We first consider general 2-edge-connected topologies, and then turn to unoriented rings.
3.1 Overview of Theorem˜1
We present a technical overview for the proof of Theorem˜1. Our algorithm aims to elect the node $r$ with the smallest identifier as the leader. It operates in two phases. In the first phase, each node maintains a counter that is incremented in an almost synchronized manner using the $\alpha$-synchronizer of Awerbuch [Awe85]. A node $v$ declares itself the leader once its counter reaches $\mathsf{ID}(v)$. This first phase alone yields a simple stabilizing leader election. To ensure quiescent termination, the second phase is executed, in which the elected leader announces its leadership through a DFS traversal.
3.1.1 Synchronized Counting
@@
To achieve an almost synchronized counting, we employ the $\alpha$-synchronizer of [Awe85], as follows. Each node maintains a counter variable $\mathsf{Count}(v)$, which is initialized to 0 at the start of the algorithm. A node increments its counter by one after receiving a new pulse from each of its neighbors. Formally, the counter value of a node $v$ is given by
@@
After electing itself as the leader, node $v$ freezes its counter at $\mathsf{Count}(v)=\mathsf{ID}(v)$ and does not increment it further. This action caps the counter value of any neighbor of $v$ at most $\mathsf{ID}(v)+1$. More generally, in a graph of $n\leq N$ nodes, the counter value of any node is upper bounded by $\mathsf{ID}(v)+N-1$, since any two nodes are connected by a path of length at most N1N-1.
@@ 3.1.2 DFS Notification
Intuitively, the leader $r$ is able to notify a neighbor $u$ because the counter value of node $u$ is constrained by a chain $P$, whose other end is anchored at the leader $r$, whose counter has already been frozen. To extend this notification process to the entire network, we perform a DFS, using the same notification procedure to explore each new edge. When a node finishes exploring all its remaining incident edges, it can notify its parent by sending pulses until a total of $\mathsf{ID}(r)+N+2$ pulses have been sent through this port since the start of the algorithm. When a node $u$ explores a back edge $\{u,v\}$, the node $v$ replies using the same method. When a node decides which new incident edge to explore, it prioritizes the edge with the smallest port number, so the resulting DFS tree is $T_{G}$, as defined in Section˜2.3.
@@
Recall from Section˜2.3 that $P_{a,b}$ is a shortest path from $a$ to $b$ in the directed version $\vec{G}$ of $G$ defined in Definition˜8.
@@
Both our leader election algorithm and the simulation algorithm of Censor-Hillel, Cohen, Gelles, and Sela [CCG+23] utilize DFS, but for different purposes. They perform a DFS from a pre-selected leader to find a cycle, while we use DFS to disseminate the identity of the elected leader to all nodes and to achieve quiescent termination. Coincidentally, both approaches leverage the strongly connected orientation of 2-edge-connected graphs. In our case, the orientation $\vec{G}$ is used solely in the analysis to facilitate the chain-and-anchor argument. In their case, it guarantees the existence of a Robbins cycle, which plays a critical role in their algorithm.
3.2 Overview of Theorem˜2
Our uniform algorithm for rings is divided into phases and is based on the assumption that no two identifiers are consecutive. To ensure this, during initialization we multiply every identifier by two.
The first phase of the algorithm uses a synchronization mechanism similar to the one used for general 2-edge-connected topologies. Each node sends a pulse to each of its neighbors and waits for pulses from them. This process is repeated a number of times equal to the node’s identifier.
Notice that the chain-and-anchor argument of the previous section shows that the process with maximum $\mathsf{ID}$ cannot exit the synchronization phase before the process with minimum $\mathsf{ID}$.
Once a node terminates the synchronization phase (called the competition phase in this algorithm), it enters a solitude checking phase to test whether its neighbours are still competing. This is done by sending two pulses on port 0 and waiting to receive two pulses. If the two pulses received are in the opposite direction from the ones sent (i.e., on port $1$), the node terminates in the leader state. Otherwise, the node becomes a non-leader. A non-leader first sends two pulses on port $1$ to “balance” the number of pulses sent in each direction, and then, after receiving a pulse from each direction, enters a phase in which it simply relays pulses.
By doing so, a non-leader node behaves in such a way that it is transparent to the other nodes.
The node with maximum $\mathsf{ID}$ will be the last to leave the competition phase, with all other nodes acting as relays. Thus, its solitude check will succeed, and the leader will notify all others by sending an additional pulse on port 0. This new pulse, together with the two pulses used for the leader’s solitude check, will form a group of three pulses traveling in the same direction around the ring. This group can be detected by all non-leader processes, by doing the difference on the number of pulses received from each port, which will then terminate.
High-level proof.
Consider the behaviour of the node $v$ with the minimum identifier $\mathsf{ID_{\min}}$, and its two neighbours $u$ and $w$. Once $v$ starts its solitude checking phase, both $u$ and $w$ are still in the competition phase, waiting for pulses from $v$. In the solitude checking phase, $v$ sends two pulses on port 0 (to node $w$) and waits for pulses. Since $u$ is still in its competition phase, it will send just one pulse to $v$ on port $1$ and will wait for a response from $v$ before sending any further pulses. Therefore, the next pulse $v$ receives must come from $w$ on port 0, causing $v$ to fail the solitude check.
Node $v$ will then send two pulses on port $1$ (balancing phase). We can show that at this point, both $u$ and $w$ have received from $v$ exactly $\mathsf{ID_{\min}}+2$ pulses, in an execution indistinguishable from one on a smaller ring where $v$ does not exist and $u$ and $w$ are directly connected. Afterwards, $v$ will only relay pulses, acting as an asynchronous link between $u$ and $w$.
Our proof idea builds on the argument above: all nodes except the one with maximum $\mathsf{ID}$ will eventually reach the solitude check and fail. In doing so, they behave in a way that hides their presence from their neighbours. This allows us to inductively reason on progressively smaller rings until only the node with maximum $\mathsf{ID}$ remains.
4 Non-Uniform Leader Election for 2-Edge-Connected Topologies
@@
The algorithm for Theorem˜1 is presented in Algorithm˜1, with the DFS notification subroutine detailed in Algorithm˜2. There are five main variables used in our algorithm:
@@
For ease of analysis, several positions in Algorithm˜1 and Algorithm˜2 are annotated with labels: $\mathsf{StartDFS}(v)$ (node $v$ initiates the DFS), $\mathsf{SendExplore}_{i}(v)$ (node $v$ begins sending an explore-notification on port $i$), $\mathsf{ReceiveExplore}_{i}(v)$ (node $v$ confirms receipt of an explore-notification from port $i$), $\mathsf{SendDone}_{i}(v)$ (node $v$ begins sending a done-notification on port $i$), and $\mathsf{ReceiveDone}_{i}(v)$ (node $v$ confirms receipt of a done-notification from port $i$). These labels are treated as events in the analysis.
@@
Any pulse that is sent during Algorithm˜2 is called a DFS notification pulse. Any remaining pulse is called a synchronized counting pulse. This distinction is used only in the analysis. The nodes cannot see these labels in the algorithm.
@@
  // Forcing all $\mathsf{ID}$s to be integer multiples of NN
@@ 19       $\mathsf{Leader}\mathsf{ID}\leftarrow\lfloor\mathsf{Count}/N\rfloor\cdot N$
21       Event: $\mathsf{ReceiveExplore}_{i}(v)$
@@ 13          Event: $\mathsf{ReceiveDone}_{j}(v)$
15       else  // Incoming DFS from port $h$ via a back edge
@@ 4.1 Synchronized Counting
We begin by analyzing the synchronized counting phase of our algorithm, which corresponds to the main loop of Algorithm˜1. We aim to show that the node $r$ with the smallest identifier eventually elects itself as the leader ($\mathsf{State}=\mathsf{Leader}$) while all other nodes are still in the synchronized counting phase ($\mathsf{State}=\bot$).
Algorithm˜1 starts with $\mathsf{Count}=0$. The action $\mathsf{SendAll}(1)$ is performed upon initialization and immediately after each counter increment, so $\sigma_{i}(v)=\mathsf{Count}(v)+1$ for all $i\in[\deg(v)]$. The counter is incremented by one whenever an incoming pulse causes $\min_{j\in[\deg(v)]}\rho_{j}$ to increase, so we always have $\mathsf{Count}=\min_{j\in[\deg(v)]}\rho_{j}$. There are two ways to exit the main loop of Algorithm˜1. The first is $\mathsf{Count}=\mathsf{ID}(v)$, which causes $v$ to become a leader. The second is $\rho_{i}-\sigma_{i}>1$, which causes $v$ to become a non-leader.
Within Section˜4.1, we focus on analyzing the behavior of the nodes in the synchronized counting phase. The only fact that we need to know about the DFS notification phase is that upon exiting the main loop of Algorithm˜1, the value of $\mathsf{Count}$ is permanently fixed. The analysis of the DFS notification phase is done in Section˜4.2.
Observation 10 (Outgoing synchronized counting pulses).
@@
During the synchronized counting phase, the action $\mathsf{SendAll}(1)$ is executed during initialization and immediately after each counter increment. There is no other mechanism by which node $v$ sends synchronized counting pulses in the algorithm. Since $\mathsf{Count}(v)$ starts at zero and becomes fixed when $v$ exits the main loop of Algorithm˜1, it follows that for any node $v\in V$, at any time during the algorithm, the number of synchronized counting pulses sent on each of its ports equals $\mathsf{Count}(v)+1$. ∎
@@
Observation 11 (Counter value).
@@
Since $\mathsf{Count}(v)$ starts at zero, it suffices to show that, while $v$ is in the synchronized counting phase, the counter is incremented by one whenever an incoming pulse causes $\min_{j\in[\deg(v)]}\rho_{j}$ to increase, and that there is no other mechanism for changing the counter value. This follows from the observation that a pulse arriving on port $i$ increases $\min_{j\in[\deg(v)]}\rho_{j}$ if and only if $\rho_{i}=\min_{j\in[\deg(v)]}\rho_{j}$ immediately after the pulse arrives. This is precisely the condition for incrementing the counter in Algorithm˜1. ∎
Since the counter is frozen once a node exits the synchronized counting phase, there is a simple generalization of ˜11 to any node $v$ in any phase: $\mathsf{Count}(v)\leq\min_{j\in[\deg(v)]}\rho_{j}(v)$.
@@
Suppose the port number of $e$ is $j$ at node $u$ and $i$ at node $v$. Since the counter is frozen once a node exits the synchronized counting phase, ˜11 implies that $\mathsf{Count}(u)\leq\rho_{j}(u)$. In other words, the number of pulses from $v$ that have been delivered to $u$ is at least $\mathsf{Count}(u)$. ˜10, together with the assumption that no DFS notification pulse from $v$ has been delivered to $u$, implies that $u$ receives at most $\mathsf{Count}(v)+1$ pulses from $v$, so $\rho_{j}(u)\leq\mathsf{Count}(v)+1$. Combining the two inequalities, we obtain $\mathsf{Count}(u)\leq\mathsf{Count}(v)+1$. Similarly, we also have $\mathsf{Count}(v)\leq\mathsf{Count}(u)+1$, so $|\mathsf{Count}(u)-\mathsf{Count}(v)|\leq 1$. ∎
@@
Let $u$ be the neighbor of $v$ such that the port number of $e=\{u,v\}$ at node $v$ is $i$. While $v$ is in the synchronized counting phase, by ˜10, $\sigma_{i}(v)=\mathsf{Count}(v)+1$. If $v$ has not received any DFS notification pulse from $u$, then ˜10 and Lemma˜12 imply that the number of pulses $v$ received from $u$ is $\rho_{i}(v)\leq\mathsf{Count}(u)+1\leq\mathsf{Count}(v)+2$, so
@@
For the rest of the discussion, we prove the claim. For each neighbor $u$ of $v$ in $S$, we have $\mathsf{Count}(u)\geq c$. By ˜10, $u$ has already sent $\mathsf{Count}(u)+1\geq c+1$ pulses to $v$. If (C1) has not happened yet, for each neighbor $w$ of $v$ in $V\setminus S$, we also have $\mathsf{Count}(w)\geq c$. By ˜10, $w$ has already sent $\mathsf{Count}(w)+1\geq c+1$ pulses to $v$.
@@
If, by this point, the $\mathsf{Non\text{-}Leader}$ exit condition $\rho_{i}(v)-\sigma_{i}(v)>1$ for $v$ has not been triggered yet, by ˜11, $\mathsf{Count}(v)$ will be incremented to $c+1$, as required.
Now suppose the $\mathsf{Non\text{-}Leader}$ exit condition $\rho_{i}(v)-\sigma_{i}(v)>1$ is triggered due to some incoming pulse at some point. By Lemma˜13, the pulse must be a DFS notification pulse. If this pulse comes from a node in $V\setminus S$, then (C3) has happened. Otherwise, the pulse comes from a node in $S$, meaning that this node has already exited the synchronized counting phase, so (C2) has happened. ∎
@@
By Lemma˜13, the first node to exit the synchronized counting phase cannot do so via the $\mathsf{Non\text{-}Leader}$ exit condition. Hence it suffices to show that $r$ is the first node that exits the synchronized counting phase.
Apply Lemma˜14 with $S=V$. Since $V\setminus S=\emptyset$, (C3) is vacuously false. Moreover, since $\min_{u\in V\setminus S}\mathsf{Count}(u)=\infty$, (C1) reduces to $\min_{v\in S}\mathsf{Count}(v)=\infty$, which cannot hold: By the time the first node $v$ reaches $\mathsf{Count}(v)=\mathsf{ID}(v)$, it must have already exited the synchronized counting phase and frozen its counter. Therefore, we conclude that (C2) must eventually occur, and hence some node $v$ eventually exits the synchronized counting phase.
@@
Suppose $v\neq r$. Consider the first moment when $\mathsf{Count}(v)=\mathsf{ID}(v)$. Consider any path $P$ connecting $v$ and $r$. Since no DFS notification pulse has been delivered so far, by Lemma˜12, the counter values of any two adjacent nodes in $P$ differ by at most one, so
@@
If we simply let the elected leader $r$ halt after the leader election phase and do not execute Algorithm˜2, we obtain a simple stabilizing leader election algorithm. By Lemma˜12, for any node $v\in V\setminus\{r\}$, we have $\mathsf{Count}(v)\leq\mathsf{Count}(r)+\mathsf{length}(P)<\mathsf{ID}(r)+(N-1)<\mathsf{ID}(v)$, so the $\mathsf{Leader}$ exit condition cannot be triggered.